Solve the equation
17/m+1/3=2/5m+29/6

Answer by Boreal(15235)

17/m + 1/3=2/5m + 29/6
Multiply everything by 30m
510 + 10m=12+145 m
498=135m
m=3.69
Note my interpretation of 2/5m, which is different from (2/5)m
Check using 3.7
17/3.7+1/3=2/18.5+29/6
4.6+0.33=0.11+4.83
This is 4.93=4.94, and that is very close given that I used 3.7 and not 3.688